# Revealing The Secret Power: How Algorithms Can Influence Content Visibility on Social Media

### v2

@@ Abstract
In recent years, the opaque design and the limited public understanding of social networks’ recommendation algorithms have raised concerns about potential manipulation of information exposure. While reducing content visibility, aka shadow banning, may help limit harmful content, it can also be used to suppress dissenting voices. This prompts the need for greater transparency and a better understanding of this practice.
In this paper, we investigate the presence of visibility alterations through a large-scale quantitative analysis of two Twitter/X datasets comprising over 40 million tweets from more than 9 million users, focused on discussions surrounding the Ukraine–Russia conflict and the 2024 US Presidential Elections. We use view counts to detect patterns of reduced or inflated visibility and examine how these correlate with user opinions, social roles, and narrative framings. Our analysis shows that the algorithm systematically penalizes tweets containing links to external resources, reducing their visibility by up to a factor of eight, regardless of the ideological stance or source reliability. Rather, content visibility may be penalized or favored depending on the specific accounts producing it, as observed when comparing tweets from the Kyiv Independent and RT.com or tweets by Donald Trump and Kamala Harris. Overall, our work highlights the importance of transparency in content moderation and recommendation systems in protecting the integrity of public discourse and ensuring equitable access to online platforms.
1 Introduction
@@
Crucially, recommendation algorithms can also limit the visibility of some content, e.g., hashtags related to dangerous or inappropriate topics or information that may encourage harmful behaviors [33]. This is known as “shadow banning,” “visibility filtering,” or “reduction,” and typically aims to moderate online discussions, promote the dissemination of reliable content, and minimize the likelihood of toxic interactions to foster a healthier information ecosystem [22].
@@
Beyond anecdotal evidence, researchers have begun looking into shadow banning through technical and statistical analyses (see Section 2). Overall, prior work suggests that shadow banning is relatively widespread across social media platforms [26, 6]. However, the lack of transparency in applying these techniques raises serious concerns. This prompts the need for systematic techniques to detect and measure visibility alterations and assess whether platforms employ shadow banning to favor specific narratives or limit dissenting voices.
Technical Roadmap. With this motivation in mind, we set out to evaluate the presence of visibility alterations in the wild. To assess visibility dynamics [38], our intuition is to rely on view counts, introduced on Twitter in late 2022 and once available through the now mostly defunct API. View counts have remained a relatively underexplored metric in related work. As academic access to the API has been restricted since June 2023 [31, 19], we turn to two publicly available datasets that include view counts: 1) almost 35M tweets related to the 2024 US Presidential Election, released by Balasubramanian et al. [3], and 2) over 17M tweets related to military aid in the Ukraine-Russia war, released by Baqir et al. [4]. [^1] As Twitter rebranded to X in July 2023, tweets are now called posts. Since one of our datasets was collected before the rebranding and one after, to ease presentation, in the rest of the paper, we stick to “tweets.” Our methodology relies on data analysis and network science techniques to uncover traces of shadow banning at content, user, and network levels. Overall, we identify and address the following research questions:
  • RQ1. Does the visibility of content depend on the characteristics of the information it contains?
    Our analysis examines how various content features, including the presence of external links, political bias, factuality of the sources, and ideological stance, may influence visibility.
  • RQ2. Do users experience different levels of visibility?
    We analyze the activities of prominent accounts to detect systematic variations in individual content visibility, i.e., whether some users’ content consistently receives less exposure.
  • RQ3. Does visibility vary across different communities?
    Social media users tend to form ideologically homogeneous communities through repeated interactions; thus, we investigate whether the content generated within these communities exhibits varying levels of visibility.
@@
    Analyzing visibility variations at the community level (RQ3), using retweet and reply information, does not reveal systematic penalization of specific groups of users.
@@
Overall, our work provides a valuable data-driven methodology using view counts and data science tools for detecting alterations in content exposure across different contexts. It also further highlights the importance of accessing information about content exposure to enhance transparency in social media algorithms and promote fairness and trustworthiness in online information ecosystems.
2 Related Work
@@ 3 Dataset
2024 US Elections. The second dataset was collected by Balasubramanian et al. [3] using a custom-built scraper. [^3] The dataset is available at https://github.com/sinking8/usc-x-24-us-election. The authors curated a list of 44 keywords related to the 2024 US Elections, including ‘2024 Elections,’ ‘Donald Trump,’ ‘letsgobrandon,’ ‘ultramaga,’ etc. Following retrieval, we filter the data as discussed in Section 4.1 and obtain approximately 35 million tweets with valid view counts by 4.2 million users.
@@
4 Methods
@@ 4.1 Pre-Processing
We also perform a filtering procedure to remove potentially inconsistent records, specifically, content for which view count is unavailable and content for which the number of interactions (i.e., likes, comments, retweets, or quotes) exceeds the recorded number of views, as users must view the content before engaging with it.
4.2 Factuality and Political Bias Estimation
To test the dependence of visibility variations on content-level features like political bias and factuality (cf. RQ1 in Section 1), we rely on third-party classifications to infer the political bias and factuality of the news sources referenced in the debate. More precisely, we use a dataset obtained from Media Bias/Fact Check (MBFC), an independent fact-checking organization that categorizes news outlets based on their factuality and political bias. [^4] https://mediabiasfactcheck.com/. This includes 2,190 different news outlets, along with their domain names, political leaning, and factuality, and classifies outlets across the political spectrum, from “extreme left” to “extreme right.” Additionally, some outlets are labeled as “questionable” or “conspiracy-pseudoscience” if they frequently publish misinformation, false content, or endorse conspiracy theories.
We use the MBFC dataset to assign factuality (from ‘Very High’ to ‘Very Low’) and political bias (from ‘Extreme Right’ to ‘Extreme Left’) to tweets based on the domain referenced. For instance, a tweet with a link to a CNN article would be classified as ‘Mostly Factual’ with a ‘Left-Center’ bias, while a tweet pointing to RT News has a ‘Very Low’ factuality with a right-center bias.
4.3 Influencers Selection
@@
For the 2024 US Elections dataset, retweet interactions are not sufficiently rich to enable a reliable estimation of users’ opinions. Therefore, we use reply interactions to reconstruct the interaction network. We identify the most replied-to users and manually extract the top 100 accounts exhibiting a clear stance, either supporting or opposing Trump. This allows us to select a balanced set of influencers representing both pro- and anti-Trump positions. Using Hartigan’s dip test of unimodality [16], we confirm the bimodal nature of the latent ideological estimation distribution for both users and influencers (D=0.022621, p-value $<2.2e-16$ for users and D = 0.10646, p-value $<2.2e-16$ for influencers). This result highlights the polarized nature of the debate, as well as the reliability of our inference [12].
@@
Figure 1: P-score distributions of original content for the Ukraine-Russia War and the 2024 US Elections datasets. The color represent the different categories of tweets.
4.4 Users’ Ideological Stance Estimation
4.5 Claim Detection and Tracking
First, we use the claim extraction pipeline from Lambretta [29] to extract claims from tweets that make an objective claim and are candidates for being tracked through downstream keyword extraction or semantic search methodologies. These claims are structurally similar to those manually curated by fact-checking organizations like Snopes or PolitiFact. [^5] See www.snopes.com, https://www.politifact.com/. We sample 1,000 unique tweets sorted by several different criteria of engagement: 1) Retweet count, 2) Reply count, 3) Like count, 4) Quote count, and 5) Impression count. This sampling gives us a set of tweets that have a higher likelihood of containing a claim (and thus garnered the level of interaction on Twitter), rather than sampling tweets randomly. We then use ClaimSpotter [17] to identify tweets that contain a claim: ClaimSpotter returns a score between 0 and 1, representing how “check-worthy” an input text is. Consistent with [29], we select 0.5 as the threshold for classifying tweets containing a claim. For the filtered tweets containing a claim, we use Lambretta’s keyword identification component to extract the best set of keywords representing the claim. Finally, we retrieve similar tweets that make the same claim as the seed set using the extracted set of keywords and use these subsets of tweets to analyze different visibility patterns.
4.6 P-score
[…]
@@ 4.6 P-score
5 Detecting Visibility Variations
In this section, we set out to answer our three research questions (cf. Section 1) related to studying the occurrence of shadow banning at content, user, and network levels.
5.1 RQ1: Content Level
@@
We classify tweets based on the domain they refer to using the MBFC dataset (see Section 4.2), dividing them into five categories: “News Outlets,” “No Domain,” “Other,” “Other Social,” and “Twitter.” These refer, respectively, to tweets that contain a link to a news outlet, no link, a link to an unclassified website, a link to another social platform, and quoting a tweet.
@@
In Figure 2, we present the distribution of p-scores across tweet categories for both datasets. In both cases, the median p-score (marked by the red line) for the “No Domain” category is higher than for all others, confirming observations from Figure 1. Specifically, tweets from news outlets exhibit the lowest median p-scores (0.033 for Ukraine-Russia, 0.012 for US Elections), while “No Domain” content shows significantly higher medians (Ukraine-Russia: 0.246; US Elections: 0.056). This further supports the hypothesis that Twitter’s algorithm favors content that does not contain external URLs. Interestingly, tweets linking to unclassified websites have median p-scores close to those of news outlets (Ukraine-Russia: 0.036; US Elections: 0.014), whereas tweets quoting other tweets or linking to other social platforms are slightly less penalized (median p-scores of 0.097 and 0.150 for Ukraine-Russia and 0.029 and 0.024 for US Elections). To validate these observations, we perform a Mann-Whitney U test between all pairs of content categories to assess whether the values in one group tend to be greater than those in another [27]. The Mann-Whitney U test does not assume normality and instead evaluates whether there is a statistically significant shift between distributions. In our case, a low p-value indicates strong evidence that the values from one group are systematically higher than those from another. For both the Ukraine-Russia and US Elections datasets, the tests confirm that the “No Domain” category statistically significantly dominates all others (p-value $<2.2\times 10^{-16}$ for all comparisons).
@@
Political Bias and Factuality. Next, we analyze the dependence of visibility on the political bias and factuality of the content being shared. We apply the same methodology to assess whether visibility varies based on the political bias and factuality of the sources referenced in tweets, for both the Ukraine-Russia (Figure 3(a)3(b)) and the 2024 US Elections (Figure 3(c)3(d)) debates. For both datasets, the results show a relatively contained variation in visibility across both political bias (Ukraine-Russia: min 0.019, max 0.05; US Elections: min 0.0081, max 0.015) and factuality (Ukraine-Russia: min 0.018, max 0.04; US Elections: min 0.0086, max 0.024), suggesting that the algorithm does not systematically penalize content based on these attributes. This is also confirmed via the Mann-Whitney U test, which did not find a clearly dominant category for political leaning in the Ukraine-Russia dataset (all categories exhibit p-values $>$ 0.05 in at least one comparison). As for factuality, the tests highlight that the “Mixed” and “Low” categories tend to be penalized (maximum p-values $<2.2\times 10^{-16}$ and 0.0019, respectively), but reject the hypothesis that “Very Low” content is consistently outperformed by “High” or “Very High” categories. Similarly, in the US Elections dataset, there is no evidence of clear dominance among political leanings (all categories have p-values $>$ 0.05 in at least one comparison), although the tests detect the lower performance of the “Left” category (p-value $<2.2\times 10^{-16}$). Interestingly, for factuality, “Very High” content outperforms all other categories (p-values $<2.2\times 10^{-16}$ in all comparisons), while no clear dominance emerges among the remaining classes. However, the number of tweets in the “Very High” category is considerably smaller than in others, suggesting an occasional rather than systematic consumption of highly factual content in the US Elections debate.
@@
Topics. Next, we use the textual content of tweets to investigate whether the different topics being discussed experienced varying levels of visibility. Following the claim detection and tracking methodology discussed in Section 4.5, for the Ukraine-Russia dataset, we first filter out 2,735 seed tweets making a claim. These seed tweets are further categorized into five different topics: 1) Europe/NATO Military Support to Ukraine, 2) Criticism of US Military Support to Ukraine, 3) US Military Support to Ukraine, 4) War Related Reporting and Events, and 5) Pro-Russian Support. Keyword extraction and the subsequent matching procedure return 151,697 tweets discussing claims related to the seed topics. We applied the same procedure for US elections dataset, categorizing the seeds into the following topics: 1) Anti-Trump Campaign Messaging, 2) International Conflicts & Electoral Impact, 3) Policy Debates & Issues, 4) Pro-Trump Campaign Messaging, 5) Public Reaction & Media Coverage, 6) Scandals, Conspiracies & Legal Controversies
Figure 4 shows the distribution of the p-score values for the claim-matched tweets, divided by each theme of claims. Similar to Figure 1, the distribution of p-scores follows a log-normal shape. However, in both cases the p-scores from the different themes have a largely balanced and symmetric distribution, contributing almost uniformly to the distribution density.

### v3

# Alessandro Galeazzi1, Pujan Paudel2, Mauro Conti1, Emiliano De Cristofaro3, Gianluca Stringhini2

@@ Abstract
In recent years, the opaque design and the limited public understanding of social networks’ recommendation algorithms have raised concerns about potential manipulation of information exposure. Reducing content visibility, aka shadow banning, may help limit harmful content; however, it can also be used to suppress dissenting voices. This prompts the need for greater transparency and a better understanding of this practice.
In this paper, we investigate the presence of visibility alterations through a large-scale quantitative analysis of two Twitter/X datasets comprising over 40 million tweets from more than 9 million users, focused on discussions surrounding the Ukraine–Russia conflict and the 2024 US Presidential Elections. We use view counts to detect patterns of reduced or inflated visibility and examine how these correlate with user opinions, social roles, and narrative framings. Our analysis shows that the algorithm systematically penalizes tweets containing links to external resources, reducing their visibility by up to a factor of eight, regardless of the ideological stance or source reliability. Rather, content visibility may be penalized or favored depending on the specific accounts producing it, as observed when comparing tweets from the Kyiv Independent and RT.com or tweets by Donald Trump and Kamala Harris. Overall, our work highlights the importance of transparency in content moderation and recommendation systems to protect the integrity of public discourse and ensure equitable access to online platforms.
I Introduction
@@
Crucially, recommendation algorithms can also limit the visibility of some content, e.g., hashtags related to dangerous or inappropriate topics or information that may encourage harmful behaviors [36]. This is known as “shadow banning,” “visibility alteration,” or “reduction,” and typically aims to moderate online discussions, promote the dissemination of reliable content, and minimize the likelihood of toxic interactions to foster a healthier information ecosystem [25]. In practice, shadow banning and, in general, algorithmic alteration aimed at reducing the visibility of a user/their content might occur to control platform dynamics and for a number of reasons, ranging from mitigating spam/bot activity, algorithmic cleaning (e.g., throttling low-performing, repetitive, or self-promotional content), to prioritizing user retention (e.g., by penalizing posts linking to competing platforms).
@@
Beyond anecdotal evidence, researchers have begun looking into shadow banning through technical and statistical analyses (see Section II). Overall, prior work suggests that shadow banning is relatively widespread across social media platforms [29, 7]. However, the lack of transparency in applying these techniques raises serious concerns. This prompts the need for systematic techniques to detect and measure visibility alterations and assess whether platforms employ shadow banning to favor specific narratives, limit dissenting voices, or prioritize corporate interests.
Technical Roadmap. Prior work has investigated shadow banning by analyzing search results or auditing recommendation algorithms, revealing potentially systematic penalization of specific users or content. However, these methods have not been able to determine whether content is penalized based on specific features (e.g., the presence of URLs), particularly at the debate level or across different topics.
In this paper, we aim to address this gap and evaluate the presence of visibility alterations in the wild using view counts, a largely underexplored metric in related work introduced on Twitter in 2022 and once available through the now mostly defunct API. As academic access to the API has been restricted since June 2023 [34, 22], we turn to two publicly available datasets that include view counts: 1) almost 35M tweets related to the 2024 US Presidential Election, released by Balasubramanian et al. [4], and 2) over 17M tweets related to military aid in the Ukraine-Russia war, released by Baqir et al. [5]. [^1] As Twitter rebranded to X in July 2023, tweets are now called posts. Since one of our datasets was collected before the rebranding and one after, to ease presentation, in the rest of the paper, we stick to “tweets.” Our methodology relies on data analysis and network science techniques to uncover traces of shadow banning at the content, user, and network levels, shedding light on visibility alterations. Although the inaccessibility of the platform’s recommendation algorithm prevents us from establishing an explicit causal link to algorithmic design, our findings provide reasonable evidence in support of this hypothesis. While alternative explanations are possible, we discuss why they are unlikely to hold in Section VI.
Research Questions. Overall, we identify and address the following research questions:
  • RQ1. Does the visibility of content depend on the characteristics of the information it contains? Our analysis examines how various content features, including the presence of external links, political bias, factuality of the sources, and ideological stance, may influence visibility.
  • RQ2. Do users experience different levels of visibility? We analyze the activities of prominent accounts to detect systematic variations in individual content visibility, i.e., whether some users’ content consistently receives less exposure.
  • RQ3. Does visibility vary across different communities? Social media users tend to form ideologically homogeneous communities through repeated interactions; thus, we investigate whether the content generated within these communities exhibits varying levels of visibility.
@@
    Analyzing visibility variations at the community level (RQ3), using retweet and reply information, does not reveal systematic penalization of specific groups.
@@
Overall, our work introduces a data-driven methodology using view counts and data science tools for detecting alterations in content exposure across different contexts. It also further highlights the importance of accessing information about content exposure to enhance transparency in social media algorithms and promote fairness and trustworthiness in online information ecosystems.
Disclaimer: Throughout the paper, we refer to systematic suppression, penalization, etc., in a broad sense, also to encompass selective content treatment. That is, we do not imply that suppression is necessarily ideologically motivated.
II Related Work
@@ III Dataset
2024 US Elections. The second dataset was collected by Balasubramanian et al. [4] using a custom-built scraper. [^3] The dataset is available at https://github.com/sinking8/usc-x-24-us-election. The authors curated a list of 44 keywords related to the 2024 US Elections, including ‘2024 Elections,’ ‘Donald Trump,’ ‘letsgobrandon,’ ‘ultramaga,’ etc. Following retrieval, we filter the data as discussed in Section IV-A and obtain approximately 35 million tweets with valid view counts by 3.8 million users.
@@
As we apply several transformations to the datasets throughout the different stages of our analysis (see Section IV), we provide summary statistics for the datasets used in the main steps of this study in Table I.
Dataset Transformation #Posts #Replies #Retweets #Quotes #Accounts #Links #Domains
US Elections Main Dataset – All Content 34,762,696 23,906,163 - 3,883,501 3,844,610 2,320,840 75,876
Ukraine-Russia with Valid View Count 4,306,644 2,208,374 - 252,849 1,509,877 631,960 53,975
US Elections Network & Latent - 4,457,343 - - 493,149 - -
Ukraine-Russia Ideology Estimation - - 9,939,787 - 67,926 - -
US Elections Claim & 2,451,532 - - - 652,851 - -
Ukraine-Russia Topic Tracking 1,845,285 - - - 583,803 - -
TABLE I: Summary statistics of the dataset used in the analysis. Retweets have likes, replies, and quotes counts equal to zero, and they inherit retweet and view counts from the original tweets. Therefore, they have been used only to compute ideology in the Ukraine-Russia War dataset and excluded from the rest of the analysis.
IV Methods
@@ IV-A Pre-Processing
We also perform a filtering procedure to remove potentially inconsistent records, specifically, content for which the view count is unavailable and content for which the number of interactions (i.e., likes, comments, retweets, or quotes) exceeds the recorded number of views, as users must view the content before engaging with it.
IV-B Factuality and Political Bias Estimation
To test the dependence of visibility variations on content-level features like political bias and factuality (cf. RQ1 in Section I), we rely on third-party classifications to infer the political bias and factuality of the news sources referenced in the debate. More precisely, we use a dataset obtained from Media Bias/Fact Check (MBFC), an independent fact-checking organization that categorizes news outlets based on their factuality and political bias. [^4] https://mediabiasfactcheck.com/. This includes 2,190 different news outlets, along with their domain names, political leaning, and factuality, and classifies outlets across the political spectrum, from “extreme left” to “extreme right.” Additionally, some outlets are labeled as “questionable” or “conspiracy-pseudoscience” if they frequently publish misinformation, false content, or endorse conspiracy theories.
We use the MBFC dataset to assign factuality (from ‘Very High’ to ‘Very Low’) and political bias (‘Extreme Right’ to ‘Extreme Left’) to tweets based on the domain referenced. For example, a tweet linking to a CNN article is classified as ‘Mostly Factual’ with a ‘Left-Center’ bias, and one linking to RT News has a ‘Very Low’ factuality with a right-center bias.
IV-C Influencers Selection
@@
For the 2024 US Elections dataset, retweet interactions are not sufficiently rich to enable a reliable estimation of users’ opinions, as the dataset contains only a few hundred retweets. Therefore, we use reply interactions to reconstruct the interaction network, which, although noisier, still allows us to distinguish between the two groups. We identify the most replied-to users and manually extract the top 100 accounts exhibiting a clear stance, either supporting or opposing Trump. This allows us to select a balanced set of influencers representing both pro- and anti-Trump positions. Using Hartigan’s dip test of unimodality [19], we confirm the bimodal nature of the latent ideological estimation distribution for users and influencers in both datasets (Ukraine-Russia War: D = 0.027, p-value <2.2e16<2.2e-16 for users and D = 0.074, p-value =5.078e06=5.078e-06 for influencers; US Elections: D = 0.012, p-value <2.2e16<2.2e-16 for users and D = 0.106, p-value <2.2e16<2.2e-16 for influencers). This result highlights the polarized nature of the debate, as well as the reliability of our inference [15].
IV-D Users’ Ideological Stance Estimation
IV-E Claim Detection and Tracking
First, we use the claim extraction pipeline from Lambretta [32] to extract claims from tweets that make an objective claim and are candidates for being tracked through downstream keyword extraction or semantic search methodologies. These claims are structurally similar to those manually curated by fact-checking organizations like Snopes or PolitiFact. [^5] See www.snopes.com, https://www.politifact.com/. We sample 1,000 unique tweets sorted by several different criteria of engagement: 1) Retweet count, 2) Reply count, 3) Like count, 4) Quote count, and 5) Views count. This sampling gives us a set of tweets that have a higher likelihood of containing a claim (and thus garnered the level of interaction on Twitter), rather than sampling tweets randomly. We then use ClaimSpotter [20] to identify tweets that contain a claim: ClaimSpotter returns a score between 0 and 1, representing how “check-worthy” an input text is. Consistent with [32], we select 0.5 as the threshold for classifying tweets containing a claim. For the filtered tweets containing a claim, we use Lambretta’s keyword identification component to extract the best set of keywords representing the claim. Finally, we retrieve similar tweets that make the same claim as the seed set using the extracted set of keywords and use these subsets of tweets to analyze different visibility patterns.
IV-F P-score
$\small\text{p-score}=\frac{\text{view count}}{\text{number of followers}}$ (1)
V Detecting Visibility Variations
In this section, we set out to answer our three research questions (cf. Section I) related to studying the occurrence of shadow banning at content, user, and network levels.
@@ V Detecting Visibility Variations
Figure 1: P-score distributions of original content for the Ukraine-Russia War and the 2024 US Elections datasets. The color represents the different categories of tweets.
@@
V-A RQ1: Content Level
@@
We classify tweets based on the domain they refer to using the MBFC dataset (see Section IV-B), dividing them into five categories: “News Outlets,” “No Domain,” “Other,” “Other Social,” and “Twitter.” These refer, respectively, to tweets that contain a link to a news outlet, no link, a link to an unclassified website, a link to another social platform, and quoting a tweet.
@@
Content Ukraine-Russia US Elections
News Outlet 252,046 1,564,923
No Domain 796,673 4,317,988
Other 730,078 851,246
Other Social 65,025 483,346
Twitter 1,463 24,509
TABLE II: Tweets by content type for distributions from Figure 2.
@@
In Figure 2, we present the distribution of p-scores across tweet categories for both datasets (also see Table II). In both cases, the median p-score (marked by the red line) for the “No Domain” category is higher than for all others, confirming observations from Figure 1. Specifically, tweets from news outlets exhibit the lowest median p-scores (0.033 for Ukraine-Russia, 0.012 for US Elections), while “No Domain” content shows significantly higher medians (Ukraine-Russia: 0.246; US Elections: 0.056). This further supports the hypothesis that Twitter’s algorithm favors content that does not contain external URLs. Interestingly, tweets linking to unclassified websites have median p-scores close to those of news outlets (Ukraine-Russia: 0.036; US Elections: 0.014), whereas tweets quoting other tweets or linking to other social platforms are slightly less penalized (median p-scores of 0.097 and 0.150 for Ukraine-Russia and 0.029 and 0.024 for US Elections). To validate these observations, we perform a Mann-Whitney U test between all pairs of content categories to assess whether the values in one group tend to be greater than those in another [30]. The Mann-Whitney U test does not assume normality and instead evaluates whether there is a statistically significant shift between distributions. In our case, a low p-value indicates strong evidence that the values from one group are systematically higher than those from another. For both the Ukraine-Russia and US Elections datasets, the tests confirm that the “No Domain” category statistically significantly dominates all others (p-value <2.2e16<2.2e-16 for all comparisons).
@@
Bias Ukraine-Russia US Elections
Extreme Right 929 124,547
Right 4,211 427,986
Right-Center 9,586 165,889
Least Biased 9,838 200,011
Left-Center 11,935 484,472
Left 3,204 148,036
(a) Political Bias
Factuality Ukraine-Russia US Elections
Very Low 845 46,747
Low 940 101,638
Mixed 10,485 668,796
Mostly Factual 4,068 218,133
High 19,158 513,152
Very High 2,818 3,230
(b) Factuality
TABLE III: Number of tweets by political bias and factuality for the distributions from Figure 3.
Political Bias and Factuality. Next, we analyze the dependence of visibility on the political bias and factuality of the content being shared. We apply the same methodology to assess whether visibility varies based on the political bias and factuality of the sources referenced in tweets, for both the Ukraine-Russia War (Figure 3(a)3(b)) and the 2024 US Elections (Figure 3(c)3(d)) debates (also see Table III). For both datasets, the results show a relatively contained variation in visibility across both political bias (Ukraine-Russia: min 0.019, max 0.05; US Elections: min 0.0081, max 0.015) and factuality (Ukraine-Russia: min 0.018, max 0.04; US Elections: min 0.0086, max 0.024), suggesting that the algorithm does not systematically penalize content based on these attributes. This is also confirmed via the Mann-Whitney U test, which did not find a dominant category for political leaning in the Ukraine-Russia dataset (all categories exhibit p-values >> 0.05 in at least one comparison). As for factuality, the tests highlight that the “Mixed” and “Low” categories tend to be penalized (maximum p-values <2.2e16<2.2e-16 and 0.0019, respectively), but reject the hypothesis that “Very Low” content is consistently outperformed by “High” or “Very High” categories. Similarly, in the US Elections dataset, there is no evidence of clear dominance among political leanings (all categories have p-values >> 0.05 in at least one comparison), although the tests detect the lower performance of the “Left” category (p-value <2.2e16<2.2e-16). Interestingly, for factuality, “Very High” content outperforms all other categories (p-values <2.2e16<2.2e-16 in all comparisons), while no clear dominance emerges among the remaining classes. However, the number of tweets in the “Very High” category is considerably smaller than in others (see Table III), suggesting an occasional rather than systematic consumption of highly factual content in the US Elections debate.
@@
Topics. Next, we use the textual content of tweets to investigate whether the different topics being discussed experienced varying levels of visibility. Following the claim detection and tracking methodology discussed in Section IV-E, for the Ukraine-Russia dataset, we first filter out 2,735 seed tweets making a claim. These seed tweets are further categorized into five different topics: 1) Europe/NATO Military Support to Ukraine, 2) Criticism of US Military Support to Ukraine, 3) US Military Support to Ukraine, 4) War Related Reporting and Events, and 5) Pro-Russian Support. Keyword extraction and the subsequent matching procedure return 151,697 tweets discussing claims related to the seed topics. We applied the same procedure for US elections dataset, categorizing the seeds into the following topics: 1) Anti-Trump Campaign Messaging, 2) International Conflicts & Electoral Impact, 3) Policy Debates & Issues, 4) Pro-Trump Campaign Messaging, 5) Public Reaction & Media Coverage, 6) Scandals, Conspiracies & Legal Controversies
Figure 4 shows the distribution of the p-score values for the claim-matched tweets, divided by each theme of claims. Similar to Figure 1, the distribution follows a log-normal shape. However, in both cases the p-scores from the different themes have a largely balanced and symmetric distribution, contributing almost uniformly to the distribution density.
@@
V-B RQ2: User Level
Inferring Users’ Ideology. To perform account-level analysis, we first need to infer users’ attributes, such as their stance on the debated issue. To this end, we use latent ideology estimation (see Section IV-D) to classify users’ ideological stances. Given a set of prominent accounts known as influencers (see Section IV-C), this technique allows us to infer users’ stances based on interaction information. Matrix decomposition maps users who interacted with a similar set of influencers close to each other in a real space, while users with fewer common influencers are positioned farther apart.
Due to the absence of retweet statistics for the 2024 US Elections dataset, we estimate latent ideology using retweet interactions for the Ukraine-Russia and reply interactions for the US Elections datasets. As mentioned, we do so as the latter only includes a few hundred retweets that did not have a valid view count. Although users’ estimaton results a bit noisier, the use of replies does not affect the overall ability to infer accounts’ stances as shown in Figure 6 and Figure 7, where we can still distinguish between influencers pro/against aid and pro/against Trump. Both datasets exhibit two communities positioned at opposite ends of the [1,1][-1,1] space, revealing polarized debates with bimodal opinion distributions, as also highlighted by Baqir et al. [5] in the context of the Ukraine-Russia war and by the Dip test as reported in Section IV-C.
In each debate, we categorize accounts into two stances: supporters (ideology score <0<0) or opponents of military aid (ideology score >0>0), and supporters (ideology score >0>0) or opponents of Donald Trump (ideology score <0<0). We then calculate the average p-score for users and influencers in our dataset. To ensure the robustness of our results, we only consider accounts with more than five original tweets.
(a) Ukraine-Russia War
(b) 2024 US Elections
Figure 6: Ideology analysis on the two datasets. The bar plots show the distribution of ideology scores for both users and influencers, while the box plots display the distribution of audience ideology scores, defined as retweeters in the Ukraine-Russia War dataset and repliers in the U.S. Elections dataset, for five representative influencers from each faction. A total of 147 prominent accounts are used as influencers for the Ukraine-Russia War dataset, and 100 accounts for the U.S. Elections.